OpenStudy (anonymous):

Hi, I need help with the steps to finding the domain of: f(x)=ln(√(x-4)-3)

OpenStudy (anonymous):

Remember that anything in the argument of a square root much be non-negative (greater than or equal to 0). In this case, we have: \[ x - 4 \ge 0 \rightarrow x \ge 4. \]Also, any argument inside an ln has to be positive. In this case, we have: \[ \sqrt {x - 4} - 3 > 0. \]Can you do the rest?